OpenStudy (anonymous):

Use the product rule to simplify the expression. 2^4x2^9x2^7

OpenStudy (anonymous):

Is the answer just 2^20? Or do I have the multiply 2x2x2 and then add the exponents?

OpenStudy (anonymous):

Since the base of exponentials multiplied are the same, you'll just have to add the exponents: 2^(4x+9x+7) = 2^(13x+7)

OpenStudy (anonymous):

Okay, so it would be 2^20. Thanks!

OpenStudy (anonymous):

You're welcome

OpenStudy (anonymous):

\[2^4*2^9*2^7=2^{(4+9+7)}=2^{20}\]